The measures of the angles in a triangle are in the extended ratio of 2:5:8. Find the measures of the angles
from the smallest to the largest and include units in your answers.

Answer by dfrazzetto(283)

Sum of interior angles of a triangle always = 180 degrees
add the individual ratios to get a sum composite ratio:
2+5+8 = 15
180 / 15 = 12
Multiply individual ratios:
{2:5:8} * 12 = 24:60:96
Check:
24 + 60 + 96 = 180